A sample of 50 customers from a large normal population has a mean waiting time of 23 minutes. We know from past testing that the population standard deviation is 4 minutes. Determine a 95% confidence interval for the true mean(waiting time) of the population. Group of answer choices (34.5051, 65.4949) (21.8913, 24.1087) (23.4505, 26.5495) (44.5051, 75.4949)
Answer #1

Confidence interval for Population mean is given as below:

Confidence interval = Xbar ± Z*σ/sqrt(n)

From given data, we have

Xbar = 23

σ = 4

n = 50

Confidence level = 95%

Critical Z value = 1.96

(by using z-table)

Confidence interval = Xbar ± Z*σ/sqrt(n)

Confidence interval = 23 ± 1.96*4/sqrt(50)

Confidence interval = 23 ± 1.1087

Lower limit = 23 - 1.1087 = 21.8913

Upper limit = 23 + 1.1087 = 24.1087

Confidence interval = (21.8913, 24.1087)